# Development of a time correction algorithm for a precise synchronization of a free-running Rubidium atomic clock with the GPS Time

### v1

@@ Abstract
We present results of our study devoted to the development of a time correction algorithm needed to precisely synchronize a free-running Rubidium atomic clock with the Coordinated Universal Time (UTC). This R&D is performed in view of the Hyper-Kamiokande (HK) experiment currently under construction in Japan, which requires a synchronization with UTC and between its different experimental sites with a precision better than $100$ ns. We use a Global Navigation Satellite System (GNSS) receiver to compare a PPS and a $10$ MHz signal, generated by a free-running Rubidium clock, to the Global Positioning System (GPS) Time signal. We use these comparisons to correct the time series (time stamps) provided by the Rubidium clock signal. We fit the difference between Rubidium and GPS Time with polynomial functions of time over a certain integration time window to extract a correction of the Rubidium time stamps in offline or online mode. In online mode, the latest fit results are used for the correction until a new comparison to the GPS Time becomes available. We show that with an integration time window of around $10^{4}$ seconds, we can correct the time stamps drift caused by the frequency random walk noise of the free running Rubidium clock so that the time difference with respect to the GPS Time stay within a ±$\pm 5$ ns range in both offline or online correction mode.
@@ 1 Introduction
A precise synchronization with the Coordinated Universal Time (UTC) or with another signal is a necessity in many applications, particularly in long-baseline physics experiments including several experimental sites. A good example is long-baseline neutrino oscillation experiments, like OPERA [1] (2006-2012), T2K [2] (from 2010) and NOvA [3] (from 2014), where a beam of neutrinos is produced and characterized in a first experimental site and detected, after several hundreds of kilometers of propagation, at another site to measure a change of the beam properties. Two next generation long-baseline neutrino experiments are being built at the moment: Hyper-Kamiokande (HK) [4] that plans to start taking data in 2027 and DUNE [5][6] that should begin sometime after 2029. These experiments require a synchronization of $100$ ns or better between the different experimental sites. Moreover, multi-messenger programs that plan to compare different components of astrophysical events [7] (e.g.: gamma-ray bursts, gravitational waves, neutrino emissions of supernovae, etc.) require a synchronization with UTC of different experiments located all over the world. For instance, to enter the SuperNova Early Warning System (SNEWS) network [8], a synchronization to UTC better than $100$ ns is required.
@@
Many long-baseline physics experiments use atomic oscillators as frequency references because of their good short term stability. Among the reference oscillators available on the market, Rubidium atomic clocks are generally chosen for their affordability as it was the case for the T2K [9] and Super-Kamiokande [10] timing systems. However, Rubidium clocks usually drift away from a stable reference because of frequency drift and random walk. For synchronization to UTC, this drift usually needs to be prevented or corrected. A common solution is to discipline the average frequency of the clock to the signals of an external Global Navigation Satellite System (GNSS) receiver, with an integration time window chosen so that it does not deteriorate the short term stability of the clock. However, it presents some drawbacks like the fact that the user has little control on the setup. In case of problems (like jumps in the time comparison), it is difficult to understand where they come from (GPS Time, receiver, the master clock, etc.) and to assess the uncertainty on the synchronization to UTC. The R&D work presented in this paper and introduced in [11] is focused on designing and characterizing an alternative method that allows more freedom to the user and a better understanding of the process. It is based on known metrology techniques [12, 13]. The proposed method uses a free-running atomic clock to derive a time signal and provide time stamps. In a physics experiment these would be the time stamps of detected events. The time stamps are corrected in post-processing using comparisons of the Rubidium clock signal to GNSS Time. In that way, we can store all the information (the raw signal, the comparisons to GPS Time, the derived correction etc.) and apply the correction in either online (during the data-acquisition) or offline modes. Let us note that the GNSS time is a good approximation of the UTC, within a few nanoseconds, and it allows synchronization to UTC via a common-view technique [14]. The common-view would be performed with a national laboratory providing a local realization of UTC(k), like e.g. the NICT laboratory in Japan [15], then the conversion to UTC can be performed with the help of the Circular T of the BIPM (Bureau International des Poids et Mesures) [16] at the end of each month.
@@ 2.1 Experimental setup
Figure 1: Experimental setup used in this work. Part of the equipment is installed at the ground floor and the other part at the fifth floor. The relevant signals generated at the ground floor are transported to the fifth floor via optical fibers with the White Rabbit (WR) protocol. This particular setup mimics what could happen in underground experiments where the clock signal would be generated underground whereas the GPS antenna and receiver would be located above-ground.
The experimental setup that we used is schematized in Figure 1. It is located at the Pierre and Marie Curie (Jussieu) campus of the Sorbonne University in Paris. The setup consists of two main parts: one represents the timing generation and correction setup, that could be reproduced in the HK experiment, and the second part is related to testing the efficiency of the correction method. In the first part a Rubidium clock (Rb) in free-running mode, at the ground floor of the laboratory, generates a Pulse Per Second (PPS) signal and a $10$ MHz signal that are transported to the fifth floor with the White Rabbit (WR) protocol. The timing signals of the slave WR switch are used by a GNSS receiver as a reference for its internal clock. The receiver connected to its antenna on the roof, above the fifth floor, is used to measure time comparisons between the GPS Time and the Rubidium clock. This physical distance between the time generation part and the receiver was done on purpose to mimic what would happen in many long-baseline physics experiments. Indeed, in Hyper-Kamiokande, the Rubidium clock would be placed inside a mountain where a cavern has been dug to host the detector whereas the receiver would have to be placed outside in a valley. The second part of our experimental setup is contained in the experimental room at the ground floor and its purpose is to validate the performance of the method, it would not be reproduced in the final setup in Hyper-Kamiokande. It consists of a frequency counter measuring the frequency of the $5$ MHz signal generated by the Rubidium clock. The reference for the internal clock of the counter is an external $5$ MHz signal generated by a Passive Hydrogen Maser (PHM).
Figure 2: Overlapping Allan Standard Deviation of the Rb/PHM frequency ratio series (in blue), measured by the frequency counter before any correction, and of GPS Time vs UTC(OP) (in orange) measured by the Septentrio receiver. The main types of noises affecting the Rubidium clock stability are indicated where they are limiting the stability.
@@ 2.1.1 Rubidium clock
The Rubidium atomic clock used is the FS725 Rubidium Frequency Standard sold by Stanford Research Systems integrating a rubidium oscillator of the PRS10 model. It provides two $10$ MHz and one $5$ MHz signals with low phase white noise and its stability estimated via the Allan Standard Deviation (ASD) [17] at $1$ s is about $2\times 10^{-11}$ (see Figure 2). It also provides a PPS output with a jitter of less than $1$ ns. Its $20$ years aging was estimated to less than $5\times 10^{-9}$ and the Mean Time Before Failure is over $200,000$ hours. It can also be frequency disciplined using an external 1 PPS reference, based on GPS for instance. The FS725 is installed at the ground floor of our laboratory and its $10$ MHz and $1$ PPS output are transported to the GNSS receiver at the fifth floor.
@@ 2.1.2 White Rabbit switches
The calibration of the link allows to obtain a sub-nanosecond synchronization between switches. A White Rabbit link between two devices is characterized by specific hardware delays and fiber propagation latencies. Each WR Master and WR Slave possesses fixed transmission and reception delays ($\Delta T_{XM}$, $\Delta RXM$, $\Delta T_{XS}$, $\Delta RXS$). These delays are the cumulative result of various factors such as SFP transceiver, PCB trace, electronic component delays, and internal FPGA chip delays. Additionally, there is a reception delay on both ends caused by aligning the recovered clock signal to the inter-symbol boundaries of the data stream, referred to as the bitslide value ($\epsilon_{M}$ and $\epsilon_{S}$ in Figure 3). We can see the results of calibration process using a counter in Figure 4, the difference of PPS signals between the WR slave and master switches changes from 165 ps to 60 ps (with a 100 m long fiber). Delays introduced by the cables were subtracted to the mean values.
@@
Note that the LPNHE, as a part of the T-REFIMEVE network [20, 21], has access through a dedicated switch to the official French realization of the UTC, called UTC(OP) (for Observatoire de Paris) [22], transported from the SYRTE laboratory via White Rabbit protocol. REFIMEVE is a French national research infrastructure aiming at the dissemination of highly accurate and stable time and frequency references to more than 30 research laboratories and research infrastructures all over France. The reference signals originate from LNE-SYRTE and are mainly transported over the optical fiber backbone of RENATER, the French National Research and Education Network. The UTC(OP) signal was not used in the final experimental setup because we do not foresee to have access to such a high precision signal in HK experiment. It was however used to characterize the GPS Time signal measured by the Septentrio receiver and whose OASD is shown in Figure 2.
2.1.3 Septentrio GPS antenna and receiver
We use the Septentrio PolaNt Choke ring GNSS antenna that supports GNSS signals from many satellite constellations including GPS, GLONASS, Galileo and BeiDou. In this work, we restrict the analysis to GPS but it can easily be generalized to any subset of constellations. The antenna position has been previously measured to a precision better than 6 mm by trilateration with the help of a web-based service provided by Canadian government [23]. We use a Septentrio PolaRx5 GNSS reference receiver as a timing receiver to compare GPS Time to the Rubidium clock. The receiver performs measurements based on the $10$ MHz reference signal coming via White Rabbit from the Rubidium clock. The Rubidium clock $1$ PPS signal is also transported to the receiver via White Rabbit to allow, at initialization, to identify the $10$ MHz cycle. Note that this $1$ PPS input is kept during the whole data-taking to avoid possible phase jumps due to perturbations. The Septentrio receiver provides one measurement every $16$ min which is the middle point of the linear function fitted from the $13$ min of data from the beginning of this $16$ min time window. The results of the measurements are registered using the CGGTTS file format [24].
@@ 2.1.3 Septentrio GPS antenna and receiver
2.1.4 Passive Hydrogen Maser
A Passive Hydrogen Maser (PHM) from T4 Science was also acquired. Note that this instrument is not available anymore. This atomic clock is approximately $10$ times more expensive than a Rubidium clock but is also much more stable. Indeed, the Allan Standard Deviation (ASD), measured with our PHM in April 2022, was only of $\sim 3\times 10^{-13}$ at $1$ s and of $1.5\times 10^{-15}$ at $1$ day. The PHM provides a $1$ PPS signal as well as two outputs of $5$ MHz, two outputs of $10$ MHz, one output of $100$ MHz and a sine output of $1$ MHz as well as a $2.048$ MHz square signal. Here, we use the PHM to generate a “perfect signal" to compare our Rubidium clock to.
2.1.5 Frequency counter
The frequency counter is the 53220A model from Keysight Technologies. It has two input channels and an input for an external frequency to use as a reference for its internal oscillator. The instrument can be used to measure the frequency of a signal input at any of the two channels. The instrument either uses directly its internal oscillator or, if specified by the user, the internal oscillator can be tuned to the external reference frequency. The external reference must be a sine wave with a frequency of $1$, $5$ or $10$ MHz. The measurement resolution depends on the gate time corresponding to the integration time window: the longer the gate time, the better the resolution. The default resolution corresponds to a 0.1 s gate time.
The frequency counter was used in continuous mode to measure the Rubidium clock $5$ MHz signal frequency simultaneously to the measurements performed by the Septentrio receiver. The external frequency reference was set to be the $5$ MHz signal of the PHM and the resolution was set to $0.01$ mHz which corresponds to a relative resolution of $2\times 10^{-11}$. This resolution is good enough to measure the ASD of the Rubidium clock at low averaging times.
@@ 2.2.1 General principle
To synchronize the Rubidium time stamps to UTC, we apply a time-dependent correction (quadratic or linear) to the time series generated by the free-running Rubidium clock $\phi_{Rb}(t)$. We model the $k^{\text{th}}$ portion of the time series ($dt_{Rb,GPS}$), defined as the difference between the free-running Rb clock and the GPS Time, as a (one or two degrees) polynomial of time
$\forall t\in[t_{k-1},t_{k}],dt_{Rb,GPS}(t)=a_{k}\cdot t^{2}+b_{k}\cdot t+c_{k}.$ (2.1)
@@
$\forall t\in[t_{k},t_{k+1}],~{}\phi_{Rb,corr}(t)=\phi_{Rb}(t)-a_{k}\times t^{2}-b_{k}\times t-c_{k}.$ (2.2)
@@
In the following, we consider two types of correction: the offline and the online corrections. The difference between the two methods is illustrated in Figure 6. The offline correction consists in using the Septentrio data from the same time-window as the Rubidium signal to extract the $a_{k}$, $b_{k}$ and $c_{k}$ coefficients. This correction is called offline because it requires the Septentrio data from up to $t_{k}+\Delta t=t_{k+1}$ to correct all the time stamps between $t_{k}$ and $t_{k+1}$ so it cannot be performed in real-time (one would need to wait a time $\Delta t$ to extract the correction coefficients for the $t_{k}$ time stamp).
@@
The online correction consists in correcting the Rubidium time stamps between $t_{k}$ and $t_{k+1}$ using Septentrio data collected before $t_{k}$. One example of online correction is illustrated in Figure 6 where overlapping windows are used. This method is called online because it can be applied in real time. In the following, we will consider the most frequent possible update of the $a_{k}$, $b_{k}$ and $c_{k}$ coefficients: they will be updated every time we receive a new data point from the Septentrio receiver (every $\delta t\approx 16$ minutes in our case). This means that we have $t_{k+1}=t_{k}+\delta t$ so that the $a_{k}$, $b_{k}$ and $c_{k}$ coefficients are extracted using Septentrio data between $t_{k}-\Delta t$ and $t_{k}$ and are used to correct the time stamps between $t_{k}$ and $t_{k}+\delta t$. In that particular case every Septentrio data point will have been used in multiple fits, the number depending on the length of the fit time window $\Delta t$.

### v2

# Precise synchronization of a free-running Rubidium atomic clock with GPS Time for applications in experimental particle physics

@@ Abstract
We present results of our study devoted to the development of a time correction algorithm needed to precisely synchronize a free-running Rubidium atomic clock with the Coordinated Universal Time (UTC). This R&D is performed in view of the Hyper-Kamiokande (HK) experiment currently under construction in Japan, which requires a synchronization with UTC and between its different experimental sites with a precision better than $100$ ns. We use a Global Navigation Satellite System (GNSS) receiver to compare a PPS and a $10$ MHz signal, generated by a free-running Rubidium clock, to the Global Positioning System (GPS) Time signal. We use these comparisons to correct the time series (time stamps) provided by the Rubidium clock signal. We fit the difference between Rubidium and GPS Time with polynomial functions of time over a certain integration time window to extract a correction of the Rubidium time stamps in offline or online mode. In online mode, the latest fit results are used for the correction until a new comparison to GPS Time becomes available. We show that with an integration time window of around $10^{4}$ seconds, we can correct the time stamps drift, caused by the frequency random walk noise and the deterministic frequency drift of the free running Rubidium clock, so that the time difference with respect to GPS Time stays within a ±$\pm 5$ ns range in both offline or online correction mode. Presented results could be of interest for other experiments in the field of neutrino physics and multi-messenger astrophysics.
keywords:
timing detectors , precise timing , atomic clock , GPS , UTC
PACS:
06.30. -k , 06.30.Ft , 07.05.Fb
MSC:
00A79 , 85-05 , 85-08
journal: NIM A
[…]
[a]organization=Laboratoire de Physique Nucléaire et des Hautes Energies (LPNHE), Sorbonne Université, CNRS/IN2P3, addressline=4 place Jussieu, city=Paris, postcode=75005, country=France
[…]
[b]organization=LNE-SYRTE, Observatoire de Paris, Université PSL, CNRS, Sorbonne Université, addressline=61 avenue de l’Observatoire, city=Paris, postcode=75014, country=France
@@ 1 Introduction
A precise synchronization with the Coordinated Universal Time (UTC) or with another signal is a necessity in many applications, particularly in long-baseline physics experiments spread over several experimental sites. A good example is long-baseline neutrino oscillation experiments, like OPERA [1] (2006-2012), T2K [2] (from 2010) and NOvA [3] (from 2014), where a beam of neutrinos is produced and characterized in a first experimental site and detected, after several hundreds of kilometers of propagation, at another site to measure a change of the beam properties. Two next generation long-baseline neutrino experiments are being built at the moment: Hyper-Kamiokande (HK) [4] that plans to start taking data in 2027 and DUNE [5, 6] that should begin sometime after 2029. These experiments require a synchronization of $100$ ns or better between the different experimental sites. Moreover, multi-messenger programs that plan to compare different components of astrophysical events [7] (e.g.: gamma-ray bursts, gravitational waves, neutrino emissions of supernovae, etc.) require a synchronization with UTC of different experiments located all over the world. For instance, to enter the SuperNova Early Warning System (SNEWS) network [8], a synchronization to UTC better than $100$ ns is required.
@@
Many long-baseline physics experiments use atomic oscillators as frequency references because of their good short term stability. Among the reference oscillators available on the market, Rubidium atomic clocks are generally chosen for their affordability as it was the case for the T2K [9] and Super-Kamiokande [10] timing systems. However, Rubidium clocks usually drift away from a stable reference because of frequency drift and random walk. For synchronization to UTC, this drift usually needs to be prevented or corrected. A common solution is to discipline the average frequency of the clock to the signals of an external Global Navigation Satellite System (GNSS) receiver, with an integration time window chosen so that it does not deteriorate the short term stability of the clock. However, it presents some drawbacks like the fact that the user has little control on the setup. In case of problems (like jumps in the time comparison), it is difficult to understand where they come from (GPS Time, receiver, the master clock, etc.) and to assess the uncertainty on the synchronization to UTC. The R&D work presented in this paper and introduced in [11] is focused on designing and characterizing an alternative method that allows more freedom to the user and a better understanding of the process. It is based on known metrology techniques [12, 13]. The proposed method uses a free-running atomic clock to derive a time signal for time-tagging the events detected by a neutrino detector for instance. The time stamps are corrected in post-processing using comparisons of the Rubidium clock signal to GNSS Time. In that way, we can store all the information (the raw signal, the comparisons to GPS Time, the derived correction etc.) and apply the correction in either online (during the data-acquisition) or offline modes. Let us note that the GNSS time is a good approximation of UTC, within a few nanoseconds, and it allows synchronization to UTC via a common-view technique [14]. The common-view would be performed with a national metrology laboratory providing a local realization of UTC called UTC(k), like e.g. the NICT laboratory in Japan [15].Then the conversion to UTC can be performed with the help of the Circular T of the BIPM (Bureau International des Poids et Mesures) [16] at the end of each month.
@@ 2.1 Experimental setup
The experimental setup that we used is schematized in Figure 1.
Figure 1: Experimental setup used in this work. Part of the equipment is installed at the ground floor and the other part at the fifth floor. The relevant signals generated at the ground floor are transported to the fifth floor via optical fibers with the White Rabbit (WR) protocol. This particular setup mimics what could happen in underground experiments where the clock signal would be generated underground whereas the GNSS antenna and receiver would be located above-ground.
It is located at the Pierre and Marie Curie (Jussieu) campus of the Sorbonne University in Paris. The setup consists of two main parts: one represents the timing generation and correction setup, that could be reproduced in the HK experiment, and the second part is related to testing the efficiency of the correction method. In the first part a Rubidium clock (Rb) in free-running mode, at the ground floor of the laboratory, generates a Pulse Per Second (PPS) signal and a $10$ MHz signal that are transported to the fifth floor with the White Rabbit (WR) protocol [17]. The timing signals of the slave WR switch are used by a GNSS receiver as a reference for its internal clock. The receiver connected to its antenna on the roof, above the fifth floor, is used to measure time comparisons between GPS Time and the Rubidium clock. This physical distance between the time generation part and the receiver was done on purpose to mimic what would happen in many long-baseline physics experiments. Indeed, in Hyper-Kamiokande, the Rubidium clock would be placed inside a mountain, where a cavern has been dug to host the detector, whereas the receiver would have to be placed outside in a valley. The second part of our experimental setup is contained in the experimental room at the ground floor and its purpose is to validate the performance of the method and would thus not be reproduced in the final setup in Hyper-Kamiokande. It consists of a counter measuring the time difference between the Rubidium clock PPS signal and the French realization of UTC (called UTC(OP) for Observatoire de Paris). The UTC(OP) reference signals are available at the laboratory that is connected to the T-REFIMEVE network [18, 19] via a third WR switch.
@@ 2.1.1 Rubidium clock
The Rubidium atomic clock used is the FS725 Rubidium Frequency Standard sold by Stanford Research Systems integrating a rubidium oscillator of the PRS10 model. It provides two $10$ MHz and one $5$ MHz signals with low phase white noise and its stability estimated via the Allan Standard Deviation (ASD) [20] at $1$ s is about $2\times 10^{-11}$. It also provides a PPS output with a jitter of less than $1$ ns. Its $20$ years aging was estimated to less than $5\times 10^{-9}$ and the Mean Time Before Failure is over $200,000$ hours. In this work we use the Rubidium clock in free-running mode but it can also be frequency disciplined using an external 1 PPS reference, based on GPS for instance. The FS725 is installed at the ground floor of our laboratory and its $10$ MHz and $1$ PPS outputs are transported to the GNSS receiver at the fifth floor.
@@ 2.1.2 White Rabbit switches
The calibration of the link allows to obtain a sub-nanosecond synchronization between switches. A White Rabbit link between two devices is characterized by specific hardware delays and fiber propagation latencies. Each WR Master and WR Slave possesses fixed transmission and reception delays ($\Delta_{TXM}$, $\Delta_{RXM}$, $\Delta_{TXS}$, $\Delta_{RXS}$). These delays are the cumulative result of various factors such as SFP transceiver, PCB trace, electronic component delays, and internal FPGA chip delays. Additionally, there is a reception delay on both ends caused by aligning the recovered clock signal to the inter-symbol boundaries of the data stream, referred to as the bitslide value ($\epsilon_{M}$ and $\epsilon_{S}$ in Figure 2). We can see the results of calibration process using a counter in Figure 3, the difference of PPS signals between the WR slave and master switches changes from 165 ps to 60 ps (with a 100 m long fiber). Delays introduced by the cables were subtracted to the mean values.
@@
As a part of the T-REFIMEVE network [18, 19], the LPNHE has access through a dedicated switch to the official French realization of the UTC, called UTC(OP) (for Observatoire de Paris) [22], transported from the SYRTE laboratory via White Rabbit protocol. REFIMEVE is a French national research infrastructure aiming at the dissemination of highly accurate and stable time and frequency references to more than 30 research laboratories and research infrastructures all over France. The reference signals originate from LNE-SYRTE and are mainly transported over the optical fiber backbone of RENATER, the French National Research and Education Network.
2.1.3 Counter
The counter is the 53220A model from Keysight Technologies. Here it was used to measure the time interval between the two PPS signals: the UTC(OP) PPS reference and the one generated by the free-running Rubidium clock.
2.1.4 Septentrio GNSS antenna and receiver
We use the Septentrio PolaNt Choke ring GNSS antenna that supports GNSS signals from many satellite constellations including GPS, GLONASS, Galileo and BeiDou. In this work, we restrict the analysis to GPS but it can easily be generalized to any subset of constellations. The antenna position has been previously measured to a precision better than 6 mm by trilateration with the help of a web-based service provided by the Canadian government [23]. We use a Septentrio PolaRx5 GNSS reference receiver as a timing receiver to compare GPS Time to the Rubidium clock. The receiver performs measurements based on the $10$ MHz reference signal coming via White Rabbit from the Rubidium clock. The Rubidium clock $1$ PPS signal, also transported to the receiver via White Rabbit, is used at initialization to identify the $10$ MHz cycle. Note that this $1$ PPS input is kept during the whole data-taking to avoid possible phase jumps due to perturbations. The Septentrio receiver provides one measurement every $16$ min which is the middle point of the linear function fitted from the $13$ min of data from the beginning of this $16$ min time window. The results of the measurements are registered using the CGGTTS file format [24].
@@ 2.2.1 General principle
To synchronize the Rubidium time stamps to UTC, we apply a time-dependent correction (quadratic or linear) to the time series generated by the free-running Rubidium clock $\phi_{Rb}(t)$. We model the $k^{\text{th}}$ portion of the time series ($dt_{Rb,GPS}$), defined as the difference between the free-running Rubidium clock and GPS Time, as a (one or two degrees) polynomial of time
$\forall t\in[t_{k-1},t_{k}],~{}dt_{Rb,GPS}(t)=a_{k}\cdot t^{2}+b_{k}\cdot t+c_{k}.$ (1)
@@
$\forall t\in[t_{k},t_{k+1}],~{}\phi_{Rb,corr}(t)=\phi_{Rb}(t)-a_{k}\times t^{2}-b_{k}\times t-c_{k}.$ (2)
@@
In the following, we consider two types of correction: the offline and the online corrections. The difference between the two methods is illustrated in Figure 5.
@@
The offline correction consists in using the Septentrio data from the same time-window as the Rubidium signal to extract the $a_{k}$, $b_{k}$ and $c_{k}$ coefficients. This correction is called offline because it requires the Septentrio data from up to $t_{k}+\Delta t=t_{k+1}$ to correct all the time stamps between $t_{k}$ and $t_{k+1}$ so it cannot be performed in real-time (one would need to wait a time $\Delta t$ to extract the correction coefficients for the $t_{k}$ time stamp).
The online correction consists in correcting the Rubidium time stamps between $t_{k}$ and $t_{k+1}$ using Septentrio data collected before $t_{k}$. One example of online correction is illustrated in Figure 5 where overlapping windows are used.
Figure 6: Overlapping Allan Standard Deviation of the Rb vs UTC(OP) time difference (in blue), measured by the counter, before any correction, and of GPS Time vs UTC(OP) (in orange) measured by the Septentrio receiver. The main types of noises affecting the Rubidium clock stability are indicated where they are limiting the stability.
This method is called online because it can be applied in real time. In the following, we will consider the most frequent possible update of the $a_{k}$, $b_{k}$ and $c_{k}$ coefficients: they will be updated every time we receive a new data point from the Septentrio receiver (every $\delta t\approx 16$ minutes in our case). This means that we have $t_{k+1}=t_{k}+\delta t$ so that the $a_{k}$, $b_{k}$ and $c_{k}$ coefficients are extracted using Septentrio data between $t_{k}-\Delta t$ and $t_{k}$ and are used to correct the time stamps between $t_{k}$ and $t_{k}+\delta t$. In that particular case every Septentrio data point will have been used in multiple fits, the number depending on the length of the fit time window $\Delta t$.